OpenStudy (okdutchman7):

Please help. WILL MEDAL! @mathstudent55 @jim_thompson5910 @Kainui @sleepyjess @satellite73 @zepdrix @rebeccaxhawaii

OpenStudy (okdutchman7):

\[\sin \frac{ \theta }{ 2 } -1 =0\]

OpenStudy (okdutchman7):

Solve the equation given that \[0 \le \theta < 360\]

OpenStudy (okdutchman7):

\[\sin \frac{ \theta }{ 2 } =1\]

zepdrix (zepdrix):

\[\large\rm \sin\color{orangered}{x}=1\]Different angle this time, right?

OpenStudy (okdutchman7):

Yes

OpenStudy (okdutchman7):

So I have to find the angles where sine = 1 and double those?

zepdrix (zepdrix):

Yes, and normally, you would double the 360k as well. But since we're in this interval 0 to 360, 720k isn't going to give us any new angles besides k=0.

OpenStudy (okdutchman7):

\[\theta = 180\]

zepdrix (zepdrix):

I guess this would be a nice way to generalize that "rule" or however you want to think of it: Whenever your theta is being multiplied by something, you'll need to include the +360k. When your theta is being divided by something, you won't need it.

OpenStudy (okdutchman7):

So is my answer^^ correct?

zepdrix (zepdrix):

yes, good job. sorry internet broke

OpenStudy (okdutchman7):

That's to me too.

OpenStudy (okdutchman7):

Thank You!